# Detection of Correlated Random Vectors

### v1

@@ 5.1 Lower bounds
As discussed above, in the fully correlated scenario where $k=n$ and $d\to\infty$, the lower bounds in Theorems 5 and 6 recover the sharp threshold $d^{-1}$, proved in [EH23]. When $d$ is fixed, on the other hand, our lower bounds are not tight in general. To wit, recall that it was proved in [EH23] that strong detection is impossible if $\rho<\rho^{\star}(d)$, and $\rho^{\star}(d)$ is a certain function of $d$. As it turns out, for $d>3$, we have $\rho^{\star}(d)>1/d$, and as so our results in this paper do not improve upon [EH23]. This suggests that a more delicate analysis of (190) is required. We suspect that the loose step in our proof, is use of the bound $|\mathsf{Par}{(m,\ell)}|d^{m}$ on the number of equivalence classes in (190). As proved in Lemma 4, a sharper bound is,
@@ 5.2 Upper bounds
Count test.
We prove now Theorem 7, and mention again that Theorem 3 is just the special case where $d=1$ and $k=n$. Consider the test in (112). We start by bounding the Type-I error probability. Markov’s inequality implies that
@@
Finally, we show that for (7) to hold it is suffice that $\rho^{2}=1-o((n^{2}/k)^{-4/d})$. We start by calculating $\psi_{P}$ and $\psi_{Q}$. The following result is proved in Appendix A.
@@
Sum test.
For simplicity of notation define $\mathsf{T}(\mathsf{X},\mathsf{Y})\triangleq\sum_{i,j=1}^{n}X_{i}^{T}Y_{j}$. Let us analyze first the Type-II error probability. We have
@@
$=\rho\left\|\bar{Y}_{k}\right\|_{2}^{2}+\rho\cdot\bar{Y}_{k}^{T}\bar{Y}^{(k+1)}+\sqrt{1-\rho^{2}}\bar{Z_{1}}^{T}\bar{Y}+\bar{Z_{2}}^{T}\bar{Y},$ (223)
where we have defined $\bar{Y}\triangleq\sum_{i=1}^{n}Y_{i}$, $\bar{Y}_{k}\triangleq\sum_{i=1}^{k}Y_{i}$, $\bar{Y}^{(k+1)}\triangleq\sum_{i=k+1}^{n}Y_{i}$, $\bar{Z_{1}}=\sum_{i=1}^{k}Z_{i}$, and $\bar{Z_{2}}=\sum_{i=k+1}^{n}Z_{i}$. Note that $\bar{Y}\sim$ , ${\cal N}(0_{d},n\cdot\mathbf{I}_{d})$, $\bar{Y_{k}}\sim$ , ${\cal N}(0_{d},k\cdot\mathbf{I}_{d})$, $\bar{Z_{1}}\sim$ , ${\cal N}(0_{d},k\cdot\mathbf{I}_{d})$, and finally $\bar{Z_{2}}\sim$ ${\cal N}(0_{d},(n-k)\cdot\mathbf{I}_{d})$. We also note that
$(\bar{Z}_{1},\bar{Z}_{2})\perp\!\!\!\perp(\bar{Y},\bar{Y}_{k},\bar{Y}^{k+1}),\text{ and }\bar{Y}_{k}\perp\!\!\!\perp\bar{Y}^{k+1}.$ (224)
@@
$\mathbb{E}_{{\cal H}_{1}}\left[\mathsf{T}(\mathsf{X},\mathsf{Y})^{2}\right]$ $=\rho^{2}\mathbb{E}\left[\left\|\bar{Y}_{k}\right\|_{2}^{4}\right]+\rho^{2}\mathbb{E}\left[\left(\bar{Y}_{k}^{T}\bar{Y}^{(k+1)}\right)^{2}\right]+(1-\rho^{2})\mathbb{E}\left[\left(\bar{Z_{1}}^{T}\bar{Y}\right)^{2}\right]$
$\quad+\mathbb{E}\left[\left(\bar{Z_{2}}^{T}\bar{Y}\right)^{2}\right].$ (225)
@@
$\mathbb{E}\left[\left(\bar{Z_{1}}^{T}\bar{Y}\right)^{2}\right]$ $=\mathbb{E}[\bar{Y}^{T}\bar{Z_{1}}\bar{Z_{1}}^{T}\bar{Y}]$ (226)
$=\mathbb{E}\left[\left.\mathbb{E}\left[\bar{Y}^{T}\bar{Z_{1}}\bar{Z_{1}}^{T}\bar{Y}\right|\bar{Y}\right]\right]$ (227)
$=\mathbb{E}\left[\bar{Y}^{T}\cdot\mathbb{E}\left[\left.\bar{Z_{1}}\bar{Z_{1}}^{T}\right|\bar{Y}\right]\cdot\bar{Y}\right]$ (228)
$=\mathbb{E}\left[\bar{Y}^{T}\mathbb{E}\left[\bar{Z_{1}}\bar{Z_{1}}^{T}\right]\bar{Y}\right]$ (229)
@@
A similar calculation shows that $\mathbb{E}\left[(\bar{Z_{2}}^{T}\bar{Y})^{2}\right]=(n-k)\mathbb{E}\left[\left\|\bar{Y}\right\|_{2}^{2}\right]=(n-k)nd$. Next, we have,
@@
Recall that $\tau=\delta\cdot\rho kd$, for some $\delta<1$. Then,
@@
Thus, we see that if $d\to\infty$, $\rho^{2}d\frac{k^{2}}{n^{2}}\to\infty$, $\rho^{2}d\frac{k}{n}\to\infty$, and $\frac{dk}{n}\to\infty$, then we have $\mathbb{P}_{{\cal H}_{1}}\left[\phi(\mathsf{X},\mathsf{Y})=0\right]\to 0$. Note however that $\rho^{2}d\frac{k^{2}}{n^{2}}\to\infty$ is more stringent compared to $\frac{dk}{n}\to\infty$ and $\rho^{2}d\frac{k}{n}\to\infty$; therefore, $\mathbb{P}_{{\cal H}_{1}}\left[\phi(\mathsf{X},\mathsf{Y})=0\right]\to 0$ if $d\to\infty$ and $\rho^{2}d\frac{k^{2}}{n^{2}}\to\infty$. As for the Type-I error probability, by substituting $\rho=0$ in the variance calculation, we get that
$\mathbb{P}_{{\cal H}_{0}}\left[\phi(\mathsf{X},\mathsf{Y})=1\right]\leq\frac{n^{2}}{(1-\delta)^{2}\rho^{2}k^{2}d},$ (241)
@@
Comparison test.
We now prove Theorem 9. We analyze the case where $\rho\in(0,1]$, with the understanding that the case where $\rho\in[-1,0)$ is analyzed in the same way. Let $G_{1}\triangleq\sum_{i=1}^{n}\sum_{j=1}^{d}X_{ij}$ and $G_{2}\triangleq\sum_{i=1}^{n}\sum_{j=1}^{d}Y_{ij}$. Then, under ${\cal H}_{0}$, we clearly have $G_{1}-G_{2}\sim{\cal N}(0,2nd)$, while under ${\cal H}_{1}$, we have $G_{1}-G_{2}\sim{\cal N}(0,2nd-2kd\rho))$. Therefore,
@@ 6 Conclusion and Outlook
In this paper, we analyzed the detection problem of deciding whether two given databases are correlated or not, under a Gaussian distributive model. With a particular focus on the canonical one-dimensional case, we derived thresholds at which optimal testing is information-theoretically impossible and possible. Our lower bounding technique is based on an orthogonal polynomial expansion of the likelihood function, which revealed interesting connection to integer partition functions. While the study of the database alignment is relatively new, there are many open questions going forward. We thereby mention several directions of particular interest for future research:

### v3

@@ 5.1 Lower bounds
As discussed above, in the fully correlated scenario where $k=n$ and $d\to\infty$, the lower bounds in Theorems 5 and 6 recover the sharp threshold $d^{-1}$, proved in [6]. When $d$ is fixed, on the other hand, our lower bounds are not tight in general. To wit, recall that it was proved in [6] that strong detection is impossible if $\rho<\rho^{\star}(d)$, and $\rho^{\star}(d)$ is a certain function of $d$. As it turns out, for $d\geq 4$, we have $\rho^{\star}(d)>1/d$, and so as our results in this paper do not improve upon [6]. This suggests that a more delicate analysis of (156) is required. We suspect that the loose step in our proof, is use of the bound $|\mathsf{Par}{(m,\ell)}|d^{m}$ on the number of equivalence classes in (156). As proved in Lemma 4, a sharper bound is,
@@ 5.2 Upper bounds
Count test. We prove now Theorem 7, and mention again that Theorem 3 is just the special case where $d=1$ and $k=n$. Consider the test in (102). We start by bounding the Type-I error probability. Markov’s inequality implies that
@@
Finally, we show that for (106) to hold it is suffice that $\rho^{2}=1-o((n^{2}/k)^{-4/d})$. We start by calculating $\psi_{P}$ and $\psi_{Q}$. The following result is proved in Appendix .1.
@@
Sum test. For simplicity of notation define $\mathsf{T}(\mathsf{X},\mathsf{Y})\triangleq\sum_{i,j=1}^{n}X_{i}^{T}Y_{j}$. Let us analyze first the Type-II error probability. We have
@@
$=\rho\left\|\bar{Y}_{k}\right\|_{2}^{2}+\rho\cdot\bar{Y}_{k}^{T}\bar{Y}^{(k+1)}+\sqrt{1-\rho^{2}}\bar{Z}_{k}^{T}\bar{Y}+(\bar{Z}^{k+1})^{T}\bar{Y},$ (190)
where we have defined $\bar{Y}\triangleq\sum_{i=1}^{n}Y_{i}$, $\bar{Y}_{k}\triangleq\sum_{i=1}^{k}Y_{i}$, $\bar{Y}^{(k+1)}\triangleq\sum_{i=k+1}^{n}Y_{i}$, $\bar{Z}_{k}=\sum_{i=1}^{k}Z_{i}$, and $\bar{Z}^{k+1}=\sum_{i=k+1}^{n}Z_{i}$. Note that $\bar{Y}\sim$ , ${\cal N}(0_{d},n\cdot\mathbf{I}_{d})$, $\bar{Y_{k}}\sim$ , ${\cal N}(0_{d},k\cdot\mathbf{I}_{d})$, $\bar{Z}_{k}\sim$ , ${\cal N}(0_{d},k\cdot\mathbf{I}_{d})$, and finally $\bar{Z}^{k+1}\sim$ ${\cal N}(0_{d},(n-k)\cdot\mathbf{I}_{d})$. We also note that
$(\bar{Z}_{k},\bar{Z}_{2})\perp\!\!\!\perp(\bar{Y},\bar{Y}_{k},\bar{Y}^{k+1}),\text{ and }\bar{Y}_{k}\perp\!\!\!\perp\bar{Y}^{k+1}.$ (191)
@@
$\mathbb{E}_{{\cal H}_{1}}\left[\mathsf{T}(\mathsf{X},\mathsf{Y})^{2}\right]$ $=\rho^{2}\mathbb{E}\left[\left\|\bar{Y}_{k}\right\|_{2}^{4}\right]+\rho^{2}\mathbb{E}\left[\left(\bar{Y}_{k}^{T}\bar{Y}^{(k+1)}\right)^{2}\right]+(1-\rho^{2})\mathbb{E}\left[\left(\bar{Z}_{k}^{T}\bar{Y}\right)^{2}\right]$
$\quad+\mathbb{E}\left[\left((\bar{Z}^{k+1})^{T}\bar{Y}\right)^{2}\right].$ (192)
@@
$\mathbb{E}\left[\left(\bar{Z}_{k}^{T}\bar{Y}\right)^{2}\right]$ $=\mathbb{E}[\bar{Y}^{T}\bar{Z}_{k}\bar{Z}_{k}^{T}\bar{Y}]$ (193)
$=\mathbb{E}\left[\left.\mathbb{E}\left[\bar{Y}^{T}\bar{Z}_{k}\bar{Z}_{k}^{T}\bar{Y}\right|\bar{Y}\right]\right]$ (194)
$=\mathbb{E}\left[\bar{Y}^{T}\cdot\mathbb{E}\left[\left.\bar{Z}_{k}\bar{Z}_{k}^{T}\right|\bar{Y}\right]\cdot\bar{Y}\right]$ (195)
$=\mathbb{E}\left[\bar{Y}^{T}\mathbb{E}\left[\bar{Z}_{k}\bar{Z}_{k}^{T}\right]\bar{Y}\right]$ (196)
@@
A similar calculation shows that $\mathbb{E}\left[((\bar{Z}^{k+1})^{T}\bar{Y})^{2}\right]=(n-k)\mathbb{E}\left[\left\|\bar{Y}\right\|_{2}^{2}\right]=(n-k)nd$. Next, we have,
@@
In the same way, we get that $\mathbb{E}_{{\cal H}_{0}}[\mathsf{T}(\mathsf{X},\mathsf{Y})]=0$ and $\mathsf{Var}_{{\cal H}_{0}}\left(\mathsf{T}(\mathsf{X},\mathsf{Y})\right)=(n-k)nd$. Recall that $\tau=\delta\cdot\rho kd$, for some $\delta<1$. Then,
@@
Thus, we see that if $d\to\infty$, $\rho^{2}d\frac{k^{2}}{n^{2}}\to\infty$, $\rho^{2}d\frac{k}{n}\to\infty$, and $\frac{dk}{n}\to\infty$, then we have $\mathbb{P}_{{\cal H}_{1}}\left[\phi(\mathsf{X},\mathsf{Y})=0\right]\to 0$. Note however that $\rho^{2}d\frac{k^{2}}{n^{2}}\to\infty$ is more stringent compared to $\frac{dk}{n}\to\infty$ and $\rho^{2}d\frac{k}{n}\to\infty$; therefore, $\mathbb{P}_{{\cal H}_{1}}\left[\phi(\mathsf{X},\mathsf{Y})=0\right]\to 0$ if $d\to\infty$ and $\rho^{2}d\frac{k^{2}}{n^{2}}\to\infty$. As for the Type-I error probability, we get that
$\mathbb{P}_{{\cal H}_{0}}\left[\phi(\mathsf{X},\mathsf{Y})=1\right]$ $\leq\frac{\mathsf{Var}_{{\cal H}_{0}}\left(\mathsf{T}(\mathsf{X},\mathsf{Y})\right)}{\left(\tau-\mathbb{E}_{{\cal H}_{0}}[\mathsf{T}(\mathsf{X},\mathsf{Y})]\right)^{2}}$ (208)
$=\frac{(n-k)nd}{\delta^{2}\rho^{2}k^{2}d^{2}}$ (209)
$\leq\frac{n^{2}}{\delta^{2}\rho^{2}k^{2}d},$ (210)
@@
Comparison test. We now prove Theorem 9. We analyze the case where $\rho\in(0,1]$, with the understanding that the case where $\rho\in[-1,0)$ is analyzed in the same way. Let $G_{1}\triangleq\sum_{i=1}^{n}\sum_{j=1}^{d}X_{ij}$ and $G_{2}\triangleq\sum_{i=1}^{n}\sum_{j=1}^{d}Y_{ij}$. Then, under ${\cal H}_{0}$, we clearly have $G_{1}-G_{2}\sim{\cal N}(0,2nd)$, while under ${\cal H}_{1}$, we have $G_{1}-G_{2}\sim{\cal N}(0,2nd-2kd\rho))$. Therefore,
@@ 6 Conclusion and Outlook
In this paper, we analyzed the detection problem of deciding whether two given databases are correlated or not, under a Gaussian distributive model. With a particular focus on the canonical one-dimensional case, we derived thresholds at which optimal testing is information-theoretically impossible and possible. Our lower bounding technique is based on an orthogonal polynomial expansion of the likelihood function, which revealed interesting connections to integer partition functions. While the study of the database alignment is relatively new, there are many open questions going forward. We thereby mention several directions of particular interest for future research: